OpenStudy (chrisga):

PLEASE HELP! What is the product of Seven square root eight times four square root five? Simplify your answer.

OpenStudy (chrisga):

Twenty eight radical ten Fifty six radical two Fifty six radical ten Two hundred eighty radical four

OpenStudy (anonymous):

Is this the problem \( \huge 7\sqrt{8} \times 4\sqrt{5} \) ??

OpenStudy (chrisga):

yes there both is inside ( )

OpenStudy (anonymous):

What one do you think is the correct answer?

OpenStudy (chrisga):

A

OpenStudy (anonymous):

It is C

OpenStudy (chrisga):

What did you multiply

OpenStudy (anonymous):

You multiply the 7 * 4 = 28 then you multiply \( \huge \sqrt{8} * \sqrt{5} = \sqrt{40} \) Also quick note, this symbol * means to times Now we have \( \huge 28 * \sqrt{40} \) but we need to break down the \( \huge \sqrt{40} \) and we can do that by finding a perfect square and a non perfect square and then timing them together. A perfect square would be \( \huge \sqrt{4} = 2 \) and a non perfect square would be \( \huge \sqrt{10} \) Now notice that \( \huge 28 * \sqrt{4} * \sqrt{10} = \sqrt{40} \) so we are still using \( \huge \sqrt{40}\) but we have just changed its form. So now we have \( \huge 28 * \sqrt{4} * \sqrt{10} = 2\sqrt{10} \) Now we take \( \huge 2\sqrt{10} \) and times it by 28 so we have \( \huge 28 * 2\sqrt{10} = 56\sqrt{10}\)

OpenStudy (chrisga):

oh ok